need to find the quotient of the complex numbers (-5+i)/(-7+ i)
Answer by Alan3354(69443)

need to find the quotient of the complex numbers (-5+i)/(-7+ i)
----------
That is the quotient.
You can "rationalize" it, get the i out of the denominator.
Multiply NUM and DEN by the conjugate of the DEN -7-i
(-5+i)/(-7+ i)
= (-5+i)*(-7-i)/((-7+ i)*(-7-i))
= (35 -12i +1)/(49 + 1)
= (36 - 12i)/50
= (18 - 6i)/25